LIM X--->INFINITY X(SQUARE)+1/X+1 - AX-2B=3/2

Grade:

1 Answers

Sher Mohammad IIT Delhi
10 years ago
lim x^2+1/(x+1) - ax-2b=3/2 lim(x^2+1/x+1=x)
lim x-ax-2b=3/2
lim x(1-a)-2b=-3/2 , give a=1 and b=3/4

sher mohammad
b.tech, iit delhi